OpenStudy (anonymous):

without graphing, determine whether the following two points would be on a vertical line or a horizontal line. (3, 4), (3, 8) vertical horizontal

OpenStudy (aroub):

8-4/3-3 4/0 The slope is undefined so it's vertical.

OpenStudy (mathmate):

Since the x values are identical, the two points have the same x-values, and different y-values. This implies they lie on a vertical line.

OpenStudy (anonymous):

(5, 1), (9, 1)

OpenStudy (mathmate):

The y-values are identical, so they lie on a horizontal line.

OpenStudy (anonymous):

Use the slope formula: \[m= (y _{2}-y_{1})/(x_{2}-x_{1})\] y = (8-4)/(3-3) = 5/0 which is undefined so the line is vertical. Lines with slope m = 0 are horizontal.